OpenStudy (db1234):

The total amount of weight on an airplane cannot exceed 3,725 pounds. The plane is carrying 1,400 pounds of luggage. The average weight of each passenger on the plane is 155 pounds. Set up and solve an inequality to find the number of passengers, p, the plane could carry.

OpenStudy (freeg13):

L=1,400

OpenStudy (freeg13):

3,725-1,400 =p

OpenStudy (freeg13):

2,325 pounds just for people

OpenStudy (freeg13):

so divide 155 into 2,325

OpenStudy (freeg13):

15

OpenStudy (will.h):

Okay.. The first thing we need to know is that the weight MUST be under 3725 so we've figured out the 1st side of the inequality, which is \[\ge 3725\] but what about the left side. the given information we have is that the plane has 1400 of luggage on board. and each passenger weight 155 pounds (Average) but we still don't really know the number of passengers so what we have so far is x= number of passengers 1400 pounds is the weight of luggage and 155 the average weight of passengers which forms the following inequality \[155p+1400 \ge 3725\] you can solve for p to find the number of passengers on board. Let me know if you need further explanation.

OpenStudy (freeg13):

again the answer is 15

OpenStudy (freeg13):

@Will.H

OpenStudy (will.h):

yes the answer is 15, nice work mate :)

OpenStudy (freeg13):

I just wanted to ask if you had any more questions